OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Part 1: Find the solution to the equation . (4 points) Part 2: Using complete sentences, explain the procedure used to solve this equation. (4 points)

OpenStudy (magbak):

\[\log_{16} 32 = x + 2 \] @primeralph

OpenStudy (magbak):

@campbell_st

OpenStudy (magbak):

@primeralph

OpenStudy (campbell_st):

well its simply a case of subtracting 2 from each side of the equation \[\log_{16}(32) - 2 = x\] job done

OpenStudy (campbell_st):

wow... and thats a 4 mark question...

OpenStudy (magbak):

what is a four mark question

OpenStudy (campbell_st):

find the solution... is 4 marks...

OpenStudy (magbak):

Oh

OpenStudy (magbak):

So how do we solve it PLEASE

OpenStudy (campbell_st):

its solved... you have a value for x

OpenStudy (magbak):

I have 3 more question can you still help me in those pelase

OpenStudy (magbak):

CAn you please explain

OpenStudy (campbell_st):

explain what all I did was subtract 2 from either side of the equation. Its an exact value answer

OpenStudy (magbak):

OK I am sorry I am a little fried.

OpenStudy (magbak):

Can it be simplified more than this

OpenStudy (anonymous):

nope

OpenStudy (campbell_st):

well it depends do you require an answer to several decimal places...

OpenStudy (magbak):

It asks for the solution

OpenStudy (campbell_st):

well you have a solution but you may need to apply change of base so that you can get an answer with a scientific calculator change of base uses base e logs ln or base 10 logs log \[\log_{a}(b) = \frac{\ln(b)}{\ln(a)}\] just apply this to your log

OpenStudy (magbak):

no it is good like that I will tag you in the next question